OpenStudy (anonymous):

Calculus: Find all the possible real numbers a,b for which the function > x^2+ax+1, if x<2 x^3+1, if x/= 3 is continuous on R.

OpenStudy (anonymous):

....hmmm I'm thinking the second piece should be: x^3 + 1 if 2<= x <3

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} x^2+ax+1 & \text{if} & x<2 \\ x^3+1& \text{if} & 2\leq x<3\\ 4x^2+b& \text{if} & x\geq 3\\ \end{array} \right. \] like that?

OpenStudy (anonymous):

And the point is that I don't want to have a disconnect like this: